10cm3, 0.12 mol/dm3 of sodium sulfite acidified with reagent hydrochloric acid solution. Then we add 10 cm3 10g/dm3 mass concentration of hydrogen sulfide solution. How many grams of sulfur precipitates?
1
Expert's answer
2019-10-17T07:18:43-0400

Na2SO3 + 6HCl + 2Na2S = 6NaCl + 3S + 3H2O


n(Na2SO3)= 1.20*10-3 mol

n(Na2S)= 2.94*10-3 mol


m(S)=1.2*10-3*3*32=0.1152 g